OpenStudy (anonymous):

find the sum of each of the followinh series or state that the series is divergent. a) 1+1/3+1/9+1/27+... b) 1- 2/3 + 4/9 - 8/27 +... c) 1/4 - 5/16 +25/64 - 125/256 +...

OpenStudy (anonymous):

I'm not sure how to approach them and getting the common ratio... please help...

ganeshie8 (ganeshie8):

start by finding the common ratio

ganeshie8 (ganeshie8):

common ratio = `next term` / `present term`

ganeshie8 (ganeshie8):

a) 1+1/3+1/9+1/27+... if `present term` = 1, then `next term` would be 1/3 common ratio = `1/3` / `1` = 1/3

OpenStudy (anonymous):

oh... i didn't know that you divide them to get the common ratio.

ganeshie8 (ganeshie8):

Yes, see if you can find the common ratio for second series : `b) 1- 2/3 + 4/9 - 8/27 +...`

OpenStudy (anonymous):

i'm not sure because of the plus sign and minus sign, but i got 4/9 as the answer

ganeshie8 (ganeshie8):

simply take the ratio : \[r = \dfrac{-2/3}{1} = -2/3\]

OpenStudy (anonymous):

oh... i thought i need to subtract and then divide...

ganeshie8 (ganeshie8):

keep in mind, for the common ratio you just take the ratio of `next term` and `present term`

ganeshie8 (ganeshie8):

try to find the common ratio for third sequence : `c) 1/4 - 5/16 +25/64 - 125/256 +...`

OpenStudy (anonymous):

is it 5/4?

ganeshie8 (ganeshie8):

thats almost correct! it should be -5/4

ganeshie8 (ganeshie8):

Now that we have the common ratios, lets find the infinite sum if they exist... do you know the infinite sum formula ?

OpenStudy (anonymous):

yes. it's S=a/1-r

ganeshie8 (ganeshie8):

yes thats the one!

ganeshie8 (ganeshie8):

Not all infinite sums converge, do you know the criterian for an inifnite geometric sum to converge ?

OpenStudy (anonymous):

no, can you please explain... thanks!

ganeshie8 (ganeshie8):

Easy : the infinite geometric series converges only if the common ratio is between -1 and 1

OpenStudy (anonymous):

why?

ganeshie8 (ganeshie8):

thats a good question, we can work that out by taking the limit of partial sum formul

ganeshie8 (ganeshie8):

see if this helps in deriving the criterian for convergence https://www.khanacademy.org/math/precalculus/seq_induction/infinite-geometric-series/v/infinite-geometric-series

ganeshie8 (ganeshie8):

After watching that video try to figure out which of the given series converge

ganeshie8 (ganeshie8):

just apply the infinite sum formula for the series that converge

OpenStudy (anonymous):

i think i know... a) and b) are convergent and c) is divergent!